OpenStudy (anonymous):

Have a couple of questions on a timed Math Quiz that I don't Understand, Please Help me to understand, Please be patient also. They will be in comment.

OpenStudy (anonymous):

How many solutions does the following system have? y = -2/3 x + 3 and 2x + 3y = 9 one solution no solution infinitely many solutions cannot be determined

OpenStudy (rajee_sam):

the second equation is written in the standard form. So let us change the first equation also into standard form. Can you do that?

OpenStudy (rajee_sam):

@bookworm00981

OpenStudy (anonymous):

I think. @rajee_sam

OpenStudy (rajee_sam):

Ok lets write it then

OpenStudy (anonymous):

We simplify it right. combine like terms?

OpenStudy (rajee_sam):

there are no like terms. We have to write it like Ax + By = C. Bringing x and y together and getting rid of the fraction and make it look just like the second equation

OpenStudy (anonymous):

Ok, let me try.

OpenStudy (rajee_sam):

\[y = -\frac{ 2 }{ 3 }x +3\]

OpenStudy (rajee_sam):

now rewrite it

OpenStudy (anonymous):

I just went Blank... don't leave let me think for a second.

OpenStudy (rajee_sam):

no I am here

OpenStudy (rajee_sam):

I help 2 - 3 people at a time. So Even if I go in and out I will see to it that you get this

OpenStudy (anonymous):

Haha I do that too.

OpenStudy (rajee_sam):

are you doing this?

OpenStudy (rajee_sam):

you need to focus on this. You can help others later

OpenStudy (anonymous):

Trying..

OpenStudy (anonymous):

I can do math, i just can't remember the steps.

OpenStudy (anonymous):

Ok, am I supposed to multiply by three on both sides?

OpenStudy (rajee_sam):

\[y = -\frac{ 2 }{ 3 }x + 3\]\[3y = 3 (-\frac{ 2 }{ 3 } x + 3)\]\[3y = -2x +9\]Bring the -2x to the other side. \[2x + 3y = 9\]

OpenStudy (rajee_sam):

yes

OpenStudy (anonymous):

Ok, let me give it a go.

OpenStudy (rajee_sam):

now what do you see in both equations?

OpenStudy (anonymous):

Ah drat, I cant draw on the equation like if it was a drawing...

OpenStudy (anonymous):

I see three's on both sides...???

OpenStudy (rajee_sam):

The first equation when I rewrote in standard form turned out to be 2x + 3y = 9 My second equation is 2x + 3y = 9 as well

OpenStudy (anonymous):

Ok, so did we flip the equation?

OpenStudy (rajee_sam):

you sound lost?

OpenStudy (anonymous):

yes..... T_T

OpenStudy (anonymous):

sorry lost connection for a sec.

OpenStudy (rajee_sam):

Let me start over. Focus and ditch your friend for a few minutes. She will live. We have two equations given to us. Both are linear. First equation is in Slope-intercept form and the second one is in Standard form. For solving system of equations we need to have the equations in standard form.

OpenStudy (rajee_sam):

Eqn. 1 \[y = -\frac{ 2 }{ 3 }x + 3\] Eqn. 2 \[2x + 3y = 9\]

OpenStudy (rajee_sam):

I am not going to do anything to eqn. 2 for now. Now let us focus on Eqn 1

OpenStudy (anonymous):

Eqn mean equation?

OpenStudy (rajee_sam):

yes

OpenStudy (rajee_sam):

So eqn. 1 I have to rewrite in standard form.

OpenStudy (anonymous):

Ok, let me. 2/3x+y=3?

OpenStudy (rajee_sam):

I have to get rid of the fractions too

OpenStudy (rajee_sam):

so what do I do?

OpenStudy (rajee_sam):

\[y = -\frac{ 2 }{ 3 }x + 3\]\[3y = 3 (-\frac{ 2 }{ 3 }x + 3 )\]\[3y = -2x + 9\]\[2x + 3y = 9\]

OpenStudy (anonymous):

Umm...

OpenStudy (rajee_sam):

Now your first eqn. has been rewritten as 2x + 3y = 9 Did you get that part?

OpenStudy (anonymous):

Looking at your work trying too.

OpenStudy (rajee_sam):

1st step was to get rid of the fraction in front of x

OpenStudy (rajee_sam):

-2/3

OpenStudy (rajee_sam):

multiply both sides of the equation by 3 ( The denominator)

OpenStudy (rajee_sam):

3y = 3(-2/3 x + 3)

OpenStudy (anonymous):

I thought I said that earlier? Oh well.

OpenStudy (rajee_sam):

And I said yes to that And you wandered off after that So lets move on to the next. Did you understand upto now?

OpenStudy (rajee_sam):

how I rewrote y = -2/3x + 3 as 2x + 3y = 9

OpenStudy (anonymous):

Fairly well I suppose. I excell in all other subjects, yet I am like a child before confusing Algebra.

OpenStudy (rajee_sam):

not to fret.

OpenStudy (rajee_sam):

You are doing good

OpenStudy (rajee_sam):

So now I have two equations. Eqn 1 2x + 3y = 9 Eqn . 2 2x + 3y = 9

OpenStudy (rajee_sam):

What can you say about them both?

OpenStudy (anonymous):

They are alike.

OpenStudy (rajee_sam):

They are the same

OpenStudy (rajee_sam):

So they will have infinitely many solutions

OpenStudy (anonymous):

Oh. so they are are parallel.

OpenStudy (rajee_sam):

They are infact the same line

OpenStudy (anonymous):

Thanks.

OpenStudy (rajee_sam):

yw

OpenStudy (phi):

Another way to do this problem (probably the way they expect you to solve it) is put both equations in slope-intercept form: y = -2/3 x + 3 and 2x + 3y = 9 become \[ y= -\frac{2}{3}x +3 \\ y= -\frac{2}{3}x +3 \] they are the same line, and lay on top on one another, so you get an infinite number of solutions. If they had the same slope, and different y intercepts, they would be parallel and never intersect, hence no solution if they have different slopes, the lines will meet somewhere, and there is one solution (where they meet)